A boy stands still in a weighing Machine which then read 65kg,he jumps upwards at the instant just before he jumps the machine reads 85kg.what acceleration does he acquire?


1
Expert's answer
2021-11-14T17:13:46-0500

"ma=N-mg"

where "N=85g" N


"a=(N-mg)\/m=(85-65)\/65=0.31" m/s2